Solution 9.128E Problem Statement Reconsider Prob. 9–127E. Determine the change in the rate of heat addition to the cycle wheDo problem 9.128E using variable specific heat assumptions instead of constant.

Please don't use constant specific heat assumption
Answer #1

Is diagram for the given problem 9. 127€ AT A For air (p = 1.005k/18k N=1.4 P2 = 14 psi = 0.965266 bar, Ti = 60°F = 288.7K RpNow, for ideal regenerator, t=1 T- Tg-Tg * To = Tg = 568.11K Also, for HEX Tg-To = Ts-Tn Đ To = Tu = 395.15k Now, it is givenand + - T6 T7 = 2 T6 T7) TH = To - 91 (To-Tx) = #AGAZIN T = TG = 777.6 - 0.93(77716-568.11) - 582.77K. Now, for perfect Re